A gas sample occupies 15.44L at 42.0°C. What is the pressure in kilopascals, given that there are 6.73mol of a gas in the sample?
1
Expert's answer
2020-04-13T02:26:50-0400

Use Ideal Gas Law

"PV=nRT"

"V= 15.44 dm^3(\frac{1 m^3}{1000 dm3})= 15.44\times 10^{-3}m^3"

"T(K) = 273+42=315K"


"P=\frac{nRT}{V}=\frac{6.73\times 8.314\times 315 }{15.44\times 10^{-3}}=1141533 Pa = 1141kPa = 1140kPa" (3 s.f.)